OpenStudy (anonymous):

In Simple Harmonic Motion, the equation of SHM is x=Acos(ωt+θ) but it is also written as x=Asin(ωt+δ), so what is the relation between two equations? and also what is the relation between the θ and δ?

OpenStudy (anonymous):

sin and cos functions are can be interchanged ,depending upon the valpe of theta or delta eg here you can say delta = theta + pie/2

OpenStudy (mrnood):

It is the PHASE of the motion which is affected by using sin or cos. At angle 0 sin= 0 At angle 0 cos = 1 however if sin(wt+theta) then the phase is 'shifted' by theta if theta=pi/2 sin (wt+pi/2)= cos(wt)

OpenStudy (anonymous):

please write equation of SHM in both sin and cos of angular frequency ω and amplitude A if the particle is situated at \[A/\sqrt{2}\] at t=0 and is going towards mean position?

OpenStudy (anonymous):

x = A Sin (wt + θ) t=0 ,x = A/2^.5 therefore \[1/\sqrt{2}\] =Sin (θ) θ = pi/4 or 3pi/4 now dx/dt = A Cos(wt +θ) at t=0 dx/dt that is velocity is negative (towards mean postion) so Cosθ is negative therefore θ=3pi/4 x=ASin(wt +3pi/4) or x = A Cos(wt+ 3pi/4 - pi/2)